OpenStudy (anonymous):

can anyone help me on removable discontinuities in a function? I'm confused.

OpenStudy (anonymous):

do you have a specific questions? removable means "you can remove them" by defining the function to be what it ought to be at that point. for example \[f(x)=\frac{x^2-4}{x-2}\] should really be \[f(x)=x+2\] but it is not defined at x = 2. if you define it to be 4 when x = 2 then you have "removed" the discontinuity

OpenStudy (anonymous):

(x^2+3x-18)/(x^2-9)